In the long run a company that produces and sells popcorn incurs total costs of $1,050 when output is 90 canisters and $1,200 when output is 120 canisters. The popcorn company exhibits
1
Expert's answer
2016-10-28T11:55:06-0400
If in the long run a company that produces and sells popcorn incurs total costs of $1,050 when output is 90 canisters and $1,200 when output is 120 canisters, then the popcorn company exhibits increasing return to scale, because with the increase of output average total costs decrease from ATC1 = 1050/90 = $11.67 to ATC2 = 1200/120 = $10 per unit.